OpenStudy (trojanpoem):

A quantity of ice ( 5 kg) at (0 C) was added to a quantity of water ( 1 kg) at (35 C). By neglecting the amount of heat lost. Calculate the final temperate of the mixture. Given that: Cw = 4180 J/kg.k Lf = 3.33* 10^5 j/kg

OpenStudy (michele_laino):

qualitatively, we can say that the melting of the sample of ice occurs, subsequently, the corresponding water will be heated to the final temperature, namelt the equilibrium temperature \(t_e\). So we can write the subsequent equation of energy balance: \[\huge \begin{gathered} {L_f}{M_1} + {c_{{H_2}O}}{M_1}\left( {{t_e} - {t_0}} \right) = \hfill \\ \hfill \\ = {c_{{H_2}O}}{M_{water}}\left( {{t_2} - {t_e}} \right) \hfill \\ \end{gathered} \] where \(\Large t_2=35°C\), \(\Large t_0=0°C\), and \(\Large c_{H_2O}=1J/(Kg\cdot K)\) Furthermore, \(\Large M_1\) is the mass of the initial sample of ice

OpenStudy (michele_laino):

Please solve that formula for \(\Large t_e\)

OpenStudy (trojanpoem):

@Michele_Laino , I want you to solve it totally by yourself. As I am going to discuss with you something related to the final answer.

OpenStudy (michele_laino):

I got a negative result as final temperature, maybe there is a typo among your data

OpenStudy (trojanpoem):

No, that's correct, You got -60 ?

OpenStudy (trojanpoem):

@Michele_Laino

OpenStudy (michele_laino):

yes! you are right! There is no reason to use the specific heat of ice

OpenStudy (trojanpoem):

Now explain what does -60 means.

OpenStudy (michele_laino):

I studied negative temperatures, when I have to study statistical physics, so please wait, I go to retrieve my textbook of statistical physics

OpenStudy (trojanpoem):

Okie dokie !

OpenStudy (trojanpoem):

Thanks for your valuable information, time.

OpenStudy (michele_laino):

please wait, I'm thinking on your question...

OpenStudy (trojanpoem):

@Michele_Laino , I closed as I though you would retrieve your textbook first. Whenever you find the answer just tag.